OpenStudy (anonymous):

a line having a slope of 3/4 passes through the point ( -8,4).write the equation of this line in slope intercept form.

OpenStudy (helder_edwin):

u have the slope m and a point (x0,y0) the equation of the line would be \[\large y-y_0=m(x-x_0) \] do some algebra (solve for y) and you will get the equation u r looking for.

OpenStudy (anonymous):

so how would i do it because i dont know

OpenStudy (kva1992):

y=3/4x+6+4= y=3/4x +10

OpenStudy (kva1992):

all you do is plug in the numbers to the equation

OpenStudy (anonymous):

so the top equation is the answer

OpenStudy (kva1992):

nope thats the formula

OpenStudy (kva1992):

-8 is your \[x_{0}\] and 4 is your \[y_{0}\]

OpenStudy (kva1992):

and all you do is plug it in

OpenStudy (anonymous):

Oh so what would be the equation of the line in slope intercept form ?

OpenStudy (kva1992):

y=3/4x +10 did that in my head so just double check the answer

OpenStudy (anonymous):

thats the answer ?

OpenStudy (kva1992):

should be